OpenStudy (anonymous):

Having trouble with algebra. May you please try to lead me through it? 7x+14-3x= 30

OpenStudy (flying_dutchman):

7x+14-3x = 30 7x-3x+14-14 = 30-14 4x = 16 4x/4 = 16/4 x = 4

OpenStudy (anonymous):

In algebra when you have an equation and want to solve for x or whatever, as I've said this many times before, what you do to one side you must also do to the other, so what ever you do on the left side of the = sign you must also do to the right. So you have the equation 7x+14-3x=30 and assuming you're solving for x, first thing you want to do is collect like terms, so move all the x's together and the integers with themselves. \[7x-3x=30-14\] you see I moved the 14 on the other side, by subtracting, because it's the opposite of addition, so as you can see to move something on the opposite side you must do the opposite operation. \[7x-3x=30-14 \implies 4x=16 \implies \frac{ 16 }{ 4 } \implies 4\] we divide because you see the 4 is 'stuck' on to the x, meaning multiplied and do the opposite operation (division) to move it to the right side. Hope all that made sense.